5 years ago, my brother was exactly twice the age i was then. And in eight year's time, the combined age of two of us will be 50. How old do u think i am?
Answer by mananth(16946)

let my age now be x years
.
5 years ago my age = x-5 years.
brother's age = 2(x-5) years.
..
After 8 years
my age = (x-5) + 13
brother's age = 2(x-5)+13
..
(x-5)+13 +2(x-5)+13=50
x-5 +13 +2x-10+13=50
x+2x+13-10+13-5=50
3x+11=50
-11 -11
3x= 50-11
3x=39
/3
x= 39/3
x= 13 years. my age now.
...
m.ananth@hotmail.ca